Consider the function

F(x)= xsin(π/x) 0<x≤1

0 x=0

Show that f(x) is continuous but not of bounded variation


1
Expert's answer
2022-03-07T17:10:01-0500

First, let us prove that "f" is continuous. It is continuous at any "0<x\leq 1", as for any such point we can find a neighbourhood not containing zero, and thus "x\mapsto \sin(\pi/x)" is well-defined and continuous. For "x=0" let us consider the limit "\lim_{x\to 0} x\sin(\pi/x) = 0", as "|\sin(\pi/x)|\leq 1" and "x\to 0". Therefore, "f" is continuous.

Now, let us prove that it is not of bounded variation. For this we will consider the following subdivisions:

"x_0 = 0, x_1= \frac{1}{n+1/2}, x_2 = \frac{1}{n-1/2}, ... , x_{n+1}=1" (for example, for "n=3" the subdivision is "\{0, \frac{2}{7}, \frac{2}{5}, \frac{2}{3}, 1\}" )

For this subdivision the variation is given by :

"V_n(f) = \sum_{0\leq i \leq n} |f(x_{i+1})-f(x_i)|"

For "i=0" we have "|f(x_1)-f(x_0)|=\frac{1}{n+1/2}" and for "0<i<n" we have

"|\frac{1}{n-(i-1)+1/2}\cdot (-1)^{n-i+1}-\frac{1}{n-i+1/2} \cdot (-1)^{n-i}|"

We remark that these expressions have opposite signs so we have

"|\frac{1}{n-(i-1)+1/2}\cdot (-1)^{n-i+1}-\frac{1}{n-i+1/2} \cdot (-1)^{n-i}| = \frac{1}{n-(i-1)+1/2}+\frac{1}{n-i+1/2}"

Which we can rewrite as "\frac{1}{n-i+1/2}=\frac{2}{2n-2i+1}"

And for "i=n" we have "|f(x_{n+1})-f(x_n)|=\frac{2}{3}".

From all this we deduce that the total variation is of a form (as we counted each term in the sum two times : once for "i" and once for "i+1") :

"V_n(f)=2\cdot \sum_{1\leq i \leq n}\frac{2}{2n+3-2i}=4\cdot \sum_{0\leq i\leq n}\frac{1}{2i+1}"

We see that for "n\to \infty" the variation behaves like the harmonic series, so it diverges, "V_n(f)\to \infty" when "n\to \infty". Therefore, it is not of bounded variation.